### v1

# Harmony in 10-TET: From Parallel Universes to the Desargues Configuration

Abstract.
The 10-tone equal temperament (10-TET) constitutes a distinct harmonic environment with step sizes exceeding the standard semitone, yet it has historically remained on the fringes of music theory. This paper proposes a radical shift in perspective, treating $\mathbb{Z}_{10}$ not as an imperfect imitation of 12-TET, but as an independent harmonic universe with its own rigorous logic. We classify all possible harmonic systems in 10-TET using a single structural parameter $\Delta$, representing the difference between the harmonic mediants (the major and minor “thirds”) which sum to the generator (the “fifth”).
Our analysis reveals a rich landscape of possibilities. The system defined by $\Delta=1$, which is naturally the first to consider, initially appears to suffer from a fundamental structural flaw: its harmonic space fractures into two disjoint universes. Conversely, the “Wide” system ($\Delta=5$) emerges as a geometrically superior structure, isomorphic to the Desargues Configuration $(10_{3})_{1}$ and forming a fully connected, vertex-transitive graph. Between these extremes lies the “Tritone” system ($\Delta=3$), which forms a connected Decagonal Prism. Our main result is the discovery of a topological isomorphism between the “restored” Acoustic system ($\Delta=1$) and the Tritone system ($\Delta=3$): despite disjoint intervallic definitions, both generate the Generalized Petersen graph $GP(10,1)$. These results suggest that in 10-TET, the path to harmonic symmetry lies in the geometry of the projective plane and prism graphs rather than pure acoustics.
1. Introduction
The 10-tone equal temperament (10-TET) has long remained a theoretical curiosity, overshadowed by the ubiquity of the standard 12-tone system. While the system’s larger step sizes initially deterred classical adaptation, its theoretical foundations were established by Sethares [17], who demonstrated its unique capacity to support stable neutral intervals and novel timbral consonance. Practical interest in the system has steadily grown, with several compositions written for 10-TET, including some specifically composed for a 10-TET piano by Hunt [8], Senpai [16], Sevish [18], and Hideya [7]. However, while these works explored the electronic domain or software simulations, purely acoustic implementations remained elusive until the recent construction of the first decaphonic piano by renowned jazz pianist Możdżer and his collaborators [13] (see our Appendix A for more details). Despite these advances, 10-TET still lacks a comprehensive harmonic theory comparable to the rich structural tradition of Western tonal music. The goal of this paper is to provide a fully self-contained theory of possible harmonies within the system $\mathbb{Z}_{10}$. We aim to define precise analogues of Major and Minor chords, not through arbitrary selection, but via a rigorous generative process, and to construct the resulting harmonic networks—or Tonnetze—associated with them.
To achieve this, we introduce a utilitarian algorithm applicable to any $n$-TET system. We postulate that a harmonic system is defined by a “Fifth” (step size $q$), defined by using an analogy of a specially devised $n$-TET analog of the Pythagorean system, a “Major Third” ($t$), and a “Minor Third” ($s$). These intervals must satisfy the condition where the two thirds sum to the fifth:
(1) $t+s\equiv q\pmod{n}$
The distinguishing feature of any such system is the difference between the thirds, denoted by $\Delta$:
(2) $t-s\equiv\Delta\pmod{n}$
This algorithmic approach allows us to scan through all possible values of $\Delta$ for a given $n$, exhaustively classifying every possible harmonic universe the temperament can support. While we focus here on $n=10$, this methodology provides a general blueprint for exploring harmonic topologies in any equal temperament. Crucially, this framework satisfies the correspondence principle vital to scientific theory generalization: a new broader theory must contain the established theory as a special case. Indeed, when we apply our $n$-TET harmonic proposal to the standard 12-TET ($n=12$), we recover the standard Western harmonic system, including the classical Euler-Riemann Tonnetz as one of the fundamental models.
Our investigation of the $\Delta$ parameter in 10-TET leads to three principal candidates for a viable musical system:
  1. (1)
    The “Narrow” choice ($\Delta=1$), characterized by $(t,s)=(4,3)$. This corresponds to the most acoustically consonant intervals (closest to the frequency ratios 3:2 and 5:4).
  2. (2)
    The “Tritone” choice ($\Delta=3$), characterized by $(t,s)=(5,2)$. This serves as a “middle ground” solution utilizing the tritone.
  3. (3)
    The “Wide” choice ($\Delta=5$), characterized by $(t,s)=(6,1)$. This system utilizes larger, more abstract interval steps.
A surprising and central finding of this work is the topological divergence—and subsequent unification—of these systems. We demonstrate that the acoustically motivated “Narrow” system $(4,3)$ suffers from a fundamental degeneracy: it splits into two disjoint, parallel universes (the “Blue” and “Red” sets). Full modulation is possible only via non-standard “wormhole” transitions.
In contrast, the “Wide” system $(6,1)$, which we analyze in the latter part of this paper, emerges as a geometrically perfect structure. We prove that this system generates a highly connected, vertex-transitive graph isomorphic to the famous Desargues Configuration from projective geometry.
Finally, in Sections 9 and 10, we analyze the intermediate “Tritone” system ($\Delta=3$). We find that it forms a robust Decagonal Prism topology. Most notably, we prove a formal isomorphism between this Tritone system and the “Restored” Acoustic system. This reveals that the disjoint acoustic universes, when properly connected, share the exact same topological backbone as the tritone system: the Generalized Petersen Graph $GP(10,1)$. Thus, 10-TET offers two distinct paths to symmetry: the projective geometry of the Desargues configuration ($\Delta=5$) and the prismatic topology shared by the Acoustic and Tritone systems ($\Delta=1,3$).
2. From Pythagorean tuning to its $n$-tone generalization
The concept of a Pythagorean tuning for an arbitrary $n$-TET ($n$-tone equal temperament) system was originally developed in [11]. In such a system, notes are generated by the repeated application of a rational interval (a generator $g>1$) followed by octave reduction:
(3) $g^{k}\cdot 2^{m(k)},\quad k\in\mathbb{Z},$
where the exponent $m(k)\in\mathbb{Z}$ is chosen uniquely such that the resulting interval lies within the range $[1,2)$. Selecting $n$ consecutive intervals yields a complete $n$-TET system. For such systems, the tempered index serves as a metric to measure the average deviation of these Pythagorean intervals from the ideal equal-tempered steps $2^{k/n}$.
Mathematical analysis presented in [11, 12] demonstrates that among small-step Pythagorean systems, the case of $n=10$ is distinguished. With the generator $g=13/8$, the 10-step system minimizes both the Pythagorean comma and the tempered index. This provides a precise theoretical justification for considering 10-TET as a compact, mathematically optimal, and musically viable alternative to the standard 12-TET.
Furthermore, [11] provides motivation grounded in concert practice, reporting that the system has been successfully implemented on acoustic grand pianos and received positively by audiences (see Appendix A for historical details).
These findings motivate the present study, which aims to codify the 10-TET system in detail. Specifically, we establish the definitions of major and minor chords within this framework and construct its analog of the Eulerian—or more precisely, the Euler [6]–Riemann [15]—Tonnetz.
To render this paper self-contained, we summarize here the fundamental concepts and key results from [11] that underpin our subsequent analysis.
2.1. General $n$-TET Pythagorean Systems
A Pythagorean system with $n$ steps per octave is constructed using a generator $g$ of harmonic type:
$1<g=\frac{2p-1}{2^{\ell(p)}}<2,\quad\text{with }p,\ell(p)\in\mathbb{N}.$
The scale is formed by the set of values:
$g^{k}\cdot 2^{m(k)},\quad k=0,1,\dots,n-1,$
taken modulo the octave. We say the system has $n$ steps and $u$ octaves if its comma $\delta$, defined as
$\delta=\left|\frac{g}{2^{u/n}}-1\right|,$
satisfies the condition:
$\delta\leq\frac{3/2}{2^{7/12}}-1\simeq 0.0011299.$
The upper bound imposed here corresponds to the comma of the classical Pythagorean system (12 steps in one octave, based on 7 octaves in the full cycle).
2.2. Why $n=10$ Is Distinguished
An exhaustive analysis of Pythagorean systems with $p<20$, $n\leq 30$ steps, and $u\leq 10$ octaves reveals that the choice $n=10$ yields the smallest comma. Remarkably, this comma is more than three times smaller than that of the classical 12-TET Pythagorean system.
This optimum is achieved for the triple $(p,u,n)=(7,7,10)$, which defines the generator of the 10-TET Pythagorean system as:
$g=\frac{13}{8}.$
With this generator, the resulting system approximates harmonic intervals closely while remaining structurally simple. This mathematically grounds the optimality of ten steps per octave in this framework.
2.3. Labelling of Intervals and the Generator
In the 10-TET Pythagorean system, intervals are labeled with integers 0 through $9$, ordered by increasing pitch. A crucial observation for the harmonic analysis is the label assigned to the generator itself.
In the 10-TET system, the generator $g=13/8$ corresponds to **Label 7**.
Label 0 1 2 3 4 5 6 7 8 9
Interval 11 $\frac{2197}{2048}$ $\frac{32768}{28561}$ $\frac{16}{13}$ $\frac{169}{128}$ $\frac{371293}{262114}$ $\frac{256}{169}$ $\mathbf{\frac{13}{8}}$ $\frac{28561}{16384}$ $\frac{4096}{2197}$
Table 1. Intervals of the 10-TET Pythagorean system. The generator (Label 7) is highlighted.
This labelling exhibits a striking parallel with the classical 12-TET system. If we construct the 12-TET Pythagorean intervals using the standard generator $g=3/2$ and order them by pitch (labeled 0 through 11), the generator again corresponds to **Label 7**.
Label 0 1 2 3 4 5 6 7 8 9 10 11
Interval 11 $\frac{256}{243}$ $\frac{9}{8}$ $\frac{32}{27}$ $\frac{81}{64}$ $\frac{4}{3}$ $\frac{729}{512}$ $\mathbf{\frac{3}{2}}$ $\frac{128}{81}$ $\frac{27}{16}$ $\frac{16}{9}$ $\frac{243}{128}$
Table 2. Intervals of the 12-TET Pythagorean system. Note that the generator is also Label 7.
Although this shared index ($q=7$) is a numerical coincidence arising from the specific ordering of pitches in each system, it suggests that the structural role of the "Fifth" (the generator interval) in 10-TET is functionally analogous to that in 12-TET. This observation forms the basis for our definition of chords in the following sections.
3. Generalization of Major and Minor Chords
To construct a harmonic theory for the 10-TET system, we must first establish a formal definition of "Major" and "Minor" chords in the language of modular arithmetic. We begin by rigorously defining these concepts for the classical 12-TET system, which will serve as our reference model.
3.1. Mathematical Definition of Chords
In an arbitrary $n$-TET system, steps are labeled $\mathbb{Z}_{n}=\{0,1,\dots,n-1\}$. Western harmony relies on three structural intervals:
  1. (1)
    The perfect fifth, denoted by $q$.
  2. (2)
    The major third, denoted by $t$.
  3. (3)
    The minor third, denoted by $s$.
A Major chord ($D_{r}$) with root $r\in\mathbb{Z}_{n}$ is defined as the ordered triple:
(4) $D_{r}=(r,\;r+t,\;r+q)\pmod{n}$.
A Minor chord ($M_{r}$) is structurally defined via the "inverse interval" logic. The minor third $s$ must be the complement of the major third $t$ within the fifth $q$:
(5) $s+t\equiv q\pmod{n}$.
The minor chord with root $r$ is then:
(6) $M_{r}=(r,\;r+s,\;r+q)\pmod{n}$.
4. The Reference Model: The Classical 12-TET System
Before applying this framework to the 10-TET system, we must exhaustively describe the properties of the classical 12-TET system. This will provide the necessary baseline for comparison.
4.1. Solutions for Thirds in 12-TET
In 12-TET, we are given $n=12$ and the Pythagorean generator implies $q=7$. A defining characteristic of the diatonic scale is that the major third is exactly one step larger than the minor third. We impose this constraint:
$t-s\equiv 1\pmod{12}$.
Combined with $t+s\equiv 7$, we solve for $t$:
$2t\equiv 8\pmod{12}$.
This congruence admits exactly two mathematical solutions:
  1. (1)
    $2t=8\implies t=4$. Consequently $s=3$.
  2. (2)
    $2t=20\implies t=10$. Consequently $s=9$.
From the perspective of Western harmony, the second solution ($t=10$) corresponds to a minor seventh interval, which is too wide to function as a third. Thus, classical theory exclusively utilizes the first solution, defining the Major Third as 4 semitones and the Minor Third as 3 semitones.
4.2. 12-TET Chords and the Tonnetz
Using the solution $(t=4,s=3)$ and $q=7$, we generate the 12-TET chords. In the standard chromatic scale, we map the note names to integers modulo 12, such that $C=0,C\sharp=1,D=2,\dots,B=11$.
Label 0 1 2 3 4 5 6 7 8 9 10 11
Name $C$ $C\sharp$ $D$ $D\sharp$ $E$ $F$ $F\sharp$ $G$ $G\sharp$ AA $A\sharp$ $B$
Table 3. Mapping of 12-TET note names to integer labels.
In this notation the 12-TET chords are:
All 12 Major Chords (12-TET) Name Root Chord Tones $D_{0}$ $C$, 0 {0, 4, 7} $D_{1}$ $C\sharp$, 1 {1, 5, 8} $_{2} DD$, 2 {2, 6, 9} $_{3} DD\sharp$, 3 {3, 7, 10} $D_{4}$ $E$, 4 {4, 8, 11} $D_{5}$ $F$, 5 {5, 9, 0} $D_{6}$ $F\sharp$, 6 {6, 10, 1} $D_{7}$ $G$, 7 {7, 11, 2} $D_{8}$ $G\sharp$, 8 {8, 0, 3} $D_{9}$ AA, 9 {9, 1, 4} $D_{10}$ $A\sharp$, 10 {10, 2, 5} $D_{11}$ $B$, 11 {11, 3, 6}
All 12 Minor Chords (12-TET) Name Root Chord Tones $M_{0}$ $C$, 0 {0, 3, 7} $M_{1}$ $C\sharp$, 1 {1, 4, 8} $M_{2}$ $D$, 2 {2, 5, 9} $M_{3}$ $D\sharp$, 3 {3, 6, 10} $M_{4}$ $E$, 4 {4, 7, 11} $M_{5}$ $F$, 5 {5, 8, 0} $M_{6}$ $F\sharp$, 6 {6, 9, 1} $M_{7}$ $G$, 7 {7, 10, 2} $M_{8}$ $G\sharp$, 8 {8, 11, 3} $M_{9}$ AA, 9 {9, 0, 4} $M_{10}$ $A\sharp$, 10 {10, 1, 5} $M_{11}$ $B$, 11 {11, 2, 6}
Figure 1. Complete enumeration of 12-TET chords.
4.3. 12-TET Chords and Their Tonnetz Assignments
Explanation of the assignment method:
The Euler-Riemann Tonnetz for 12-TET is constructed by connecting chords that share exactly two common tones. In this system, every major chord is adjacent to exactly three minor chords:
  • Parallel (P): $D_{r}\leftrightarrow M_{r}$ (share Root, Fifth).
  • Relative (R): $D_{r}\leftrightarrow M_{r-3}$ (share Root, Third).
  • Leading-Tone (L): $D_{r}\leftrightarrow M_{r+4}$ (share Third, Fifth).
In short, the method states that:
  • Replace exactly one note of the chord with another note to form a minor chord.
  • The resulting chord must share exactly two common tones (modulo 12) with the original chord.
Example 1: The chord $D_{0}$ (C Major) = {0, 4, 7}
  • P transformation: $D_{0}=\{0,4,7\}\to\{0,3,7\}=M_{0}$ (C minor). Common tones: {0, 7}.
  • R transformation: $D_{0}=\{0,4,7\}\to\{9,0,4\}=M_{9}$ (A minor). Common tones: {0, 4}.
  • L transformation: $D_{0}=\{0,4,7\}\to\{4,7,11\}=M_{4}$ (E minor). Common tones: {4, 7}.
Example 2: The chord $D_{7}$ (G Major) = {7, 11, 2}
  • P transformation: $D_{7}=\{7,11,2\}\to\{7,10,2\}=M_{7}$ (G minor). Common tones: {7, 2}.
  • R transformation: $D_{7}=\{7,11,2\}\to\{4,7,11\}=M_{4}$ (E minor). Common tones: {7, 11}.
  • L transformation: $D_{7}=\{7,11,2\}\to\{11,2,6\}=M_{11}$ (B minor). Common tones: {11, 2}.
Similarly for minor chords: we replace one note to form a major chord, preserving two common tones. Example 3: The chord $M_{2}$ (D minor) = {2,5,9}
  • P transformation: $M_{2}=\{2,5,9\}\to\{2,6,9\}=D_{2}$ (D Major). Common tones: {2, 9}.
  • R transformation: $M_{2}=\{2,5,9\}\to\{0,2,5\}=D_{10}$ ($A\sharp$ Major). Common tones: {2, 5}.
  • L transformation: $M_{2}=\{2,5,9\}\to\{5,9,0\}=D_{5}$ (F Major). Common tones: {5, 9}.
Example 4: The chord $M_{9}$ (A minor) = {9, 0, 4}
  • P transformation: $M_{9}=\{9,0,4\}\to\{9,1,4\}=D_{9}$ (A Major). Common tones: {9, 4}.
  • R transformation: $M_{9}=\{9,0,4\}\to\{5,9,0\}=D_{5}$ (F Major). Common tones: {9, 0}.
  • L transformation: $M_{9}=\{9,0,4\}\to\{0,4,7\}=D_{0}$ (C Major). Common tones: {0, 4}.
This leads to the following Tonnetz neighbor tables in integer notation:
[…]
Figure 2. The connectivity of the 12-TET Tonnetz in integer notation.
These relationships produce the hexagonal honeycomb structure on the plane known as the Euler – Riemann Tonnetz, composed of triangles where each vertex is a chord, and adjacent chords share two tones. This is illustrated below in Figure 3 quoted from [4].
Figure 3. The 12-TET Tonnetz.
This Tonnetz can be also visualized as a Levi bipartite graph of a $12_{3}$ configuration [4]:
Figure 4. The 12-TET Tonnetz as a Levi bipartite graph of 12 points and 12 lines, sharing the incidence relation where each line contains 3 points, and every three lines meet at precisely one point [4].
5. The 10-TET System and the Choice of Thirds
We now return to Section 3.1 to define chords in 10-TET system. We have established that for the 10-TET the generator $g=13/8$ dictates the fifth $q=7$. However, the value of the major third $t$ is a free parameter. To define a harmonic system analogous to Western diatonic harmony, we must determine the relationship between the major third ($t$) and the minor third ($s$).
We recall equation (5) defining the main characteristic of the fifth in any $n$-TET system as the sum of thirds. When applied to 10-TET, this gives:
(7) $t+s\equiv 7\pmod{10}$.
We now systematically explore possible constraints on the difference $\Delta=t-s$.
5.1. Hypothesis 1: The Diatonic Difference ($\Delta=1$)
The classical 12-TET system satisfies ts=1t-s=1. Imposing this on 10-TET yields:
$2t\equiv 8\pmod{10}$.
This admits two solutions:
  • $t=4$ (implies $s=3$). This is the solution analyzed in the remainder of this paper.
  • $t=9$ (implies $s=8$). This yields large intervals (sevenths) unsuitable for tertian harmony.
5.2. Hypothesis 2: The Whole-Tone Difference ($\Delta=2$)
One might ask if the major and minor thirds could differ by a larger step, for instance, $\Delta=2$. This leads to the system:
(8) $\begin{cases}t+s\equiv 7\\ t-s\equiv 2\end{cases}$
Summing the equations gives $2t\equiv 9\pmod{10}$. In the ring of integers modulo 10, the quantity $2t$ must always be even. Since 9 is odd, **no integer solution exists**. Thus, it is mathematically impossible to construct a Major/Minor system in 10-TET where the thirds differ by 2 steps.
Of course the same argument shows that any even number $\Delta\in\mathbb{Z}_{10}$ makes the system
$t+s\equiv 7\quad\&\quad t-s\equiv\Delta\quad\&\quad\Delta\in\mathbb{Z}_{10}\,\,\text{with}\,\,\Delta\,\,\text{even},$
incompatible. Thus, we have no solutions for all $\Delta$ even.
5.3. Hypothesis 3: The “Neutral” Difference ($\Delta=3$)
Next, we test the difference $\Delta=3$:
(9) $\begin{cases}t+s\equiv 7\\ t-s\equiv 3\end{cases}$
Summing gives $2t\equiv 10\equiv 0\pmod{10}$. The solutions are:
  • $t=0$ (implies $s=7$). Trivial solution (unison and fifth).
  • $t=5$ (implies $s=2$).
The solution $(t=5,s=2)$ is mathematically valid. However, the "Major Third" $t=5$ corresponds to exactly half the octave ($\sqrt{2}$, or the tritone). A chord constructed this way would be $C=\{0,5,7\}$. While structurally interesting, a system based on the tritone as a consonance diverges radically from the acoustic principles of tertian harmony.
5.4. Hypothesis 4: The case $\Delta=5$ i.e. The Wide System
This is perhaps the most interesting case, with a solution $(t,s)=(6,1)$, and due to its significance, it requires separate treatment. We postpone its discussion to Section 8, after we perform the full analysis of the case $\Delta=1$.
5.5. Conclusion on the Choice of Thirds
This exhaustion of possibilities gives us the way to proceed, and to analyze in detail two cases only, namely cases $(t,s)=(4,3)$ and $(t,s)=(6,1)$.
We first fully analyse the case $(t,s)=(4,3)$. This choice is the **unique** solution that:
  1. (1)
    Is mathematically solvable (unlike $\Delta$ is even).
  2. (2)
    It is **narrow** in the distance between the thirds, providing the closest mathematical analogy to the standard 12-TET-system.
Therefore first, in the next section, we proceed with the analysis of the solution $(t=4,s=3)$.
6. Analysis of the Primary Solution (4,3)
We formally define the chords for 10-TET using $t=4$ and $s=3$:
$D_{r}=\{r,r+4,r+7\}\pmod{10}$
$M_{r}=\{r,r+3,r+7\}\pmod{10}$
6.1. Explicit Enumeration and the Discovery of Identity
Let us list the chords explicitly as sets of integers to inspect their structure.
Major Chords ($D_{r}$) Name Root Set $D_{0}$ 0 {0, 4, 7} $D_{1}$ 1 {1, 5, 8} $D_{2}$ 2 {2, 6, 9} $D_{3}$ 3 {3, 7, 0} $D_{4}$ 4 {4, 8, 1} $D_{5}$ 5 {5, 9, 2} $D_{6}$ 6 {6, 0, 3} $D_{7}$ 7 {7, 1, 4} $D_{8}$ 8 {8, 2, 5} $D_{9}$ 9 {9, 3, 6}
Minor Chords ($M_{r}$) Name Root Set $M_{0}$ 0 {0, 3, 7} $M_{1}$ 1 {1, 4, 8} $M_{2}$ 2 {2, 5, 9} $M_{3}$ 3 {3, 6, 0} $M_{4}$ 4 {4, 7, 1} $M_{5}$ 5 {5, 8, 2} $M_{6}$ 6 {6, 9, 3} $M_{7}$ 7 {7, 0, 4} $M_{8}$ 8 {8, 1, 5} $M_{9}$ 9 {9, 2, 6}
Figure 5. Explicit pitch sets of 10-TET chords. Note the highlighted entries.
6.2. Theorem of Modal Degeneracy
Inspection of the tables above reveals a startling fact. Compare, for example, the major chord rooted at 0 with the minor chord rooted at 7:
  • $D_{0}=\{0,4,7\}$
  • $M_{7}=\{7,0,4\}$
These are the same set. This is not an isolated anomaly but a general property of the system.
Theorem 1 (Modal Identity).
In the 10-TET system with $(t=4,s=3)$, the set of pitch classes forming a major chord $D_{r}$ is identical to the set forming the minor chord $M_{r+7}$.
$D_{r}\equiv M_{r+7}\pmod{10}$.
Proof.
$D_{r}=\{r,r+4,r+7\}$. $M_{r+7}=\{(r+7),(r+7)+3,(r+7)+7\}=\{r+7,r+10,r+14\}\equiv\{r+7,r,r+4\}$. The sets are identical. ∎
6.3. Musical Interpretation of Modal Degeneracy
This mathematical identity has profound musical consequences for the 10-TET system.
  • Structural Uniqueness: Unlike in 12-TET, where a Major triad (e.g., C-E-G) and its relative Minor (A-C-E) are distinct objects sharing two tones, in 10-TET they are the same physical object sharing all three tones. In fact, there are only 10 unique triads in the entire $(4,3)$ solution 10-TET Universe.
  • Context-Dependent Perception: Since the interval set $\{4,3,3\}$, corresponding to a passage $0\to 4\to 7\to 0\equiv 10$, is invariant under cyclic permutation, the perceived quality of the chord depends entirely on inversion and context (specifically, which note is heard as the root or bass):
    • If the pitch $r$ is established as the root, the intervals above it are $4+3$, and the chord is heard as Major ($D_{r}$).
    • If the pitch $r+7$ is established as the root, the intervals above it are $3+4$, and the chord is heard as Minor ($M_{r+7}$).
  • The Janus-Faced Triad: Physically, the 10-TET triad is an ambiguous, "Janus-faced" harmonic object. It functions as a bridge between the major and minor modes not by changing notes (voice leading), but by changing the listener’s perspective (re-interpreting the root).
6.4. The Degenerate Tonnetz as a Levi Graph
This identity fundamentally alters the topological construction of the Tonnetz. In 12-TET, the "Relative" transformation connects a chord to a distinct neighbor ($R:D_{r}\to M_{r-3}$), creating a graph where every vertex has degree 3. In 10-TET, the "Relative" transformation ($D_{r}\to M_{r+7}$) maps the chord to itself. Consequently, the edge corresponding to $R$ becomes a loop, and the functional degree of each vertex drops to 2.
To visualize this structure globally, we construct the **Levi graph** of the configuration. We treat the set of 10 major chords as "points" and the set of 10 minor chords as "lines". An incidence (edge) exists if a major chord transforms into a minor chord via a Parallel ($P$) or Leading-Tone ($L$) motion.
Since each major chord connects exactly to two minor chords, and each minor chord connects to two major chords, the global structure forms a **Levi graph of a $(10_{2})$ configuration**.
We present this graph below by arranging all 20 chords on a single circle. To highlight the structural relationships, we arrange them in the sequence: $D_{0},M_{4},D_{1},M_{5},D_{4},M_{8},\dots$. The connections reveal the hidden topology.
$D_{0}$$M_{4}$$D_{1}$$M_{5}$$D_{4}$$M_{8}$$D_{5}$$M_{9}$$D_{8}$$M_{2}$$D_{9}$$M_{3}$$D_{2}$$M_{6}$$D_{3}$$M_{7}$$D_{6}$$M_{0}$$D_{7}$$M_{1}$
Figure 6. The full Levi graph of the 10-TET Tonnetz arranged on a circle. The blue edges correspond to the connections between "even" chords, while the red edges connect "odd" chords. The interleaving of the nodes visually demonstrates the symmetry of the system.
6.4.1. Topological Decomposition
The coloring in Figure 6 makes the fundamental property of the system immediately visible. Although constructed as a single harmonic system, the graph naturally separates into two disjoint, non-intersecting subgraphs.
The blue path traces a closed cycle involving only the chords with even roots (for $D$) and their associated minors. The red path traces a similar cycle for the odd roots. No blue edge ever meets a red edge.
Theorem 2 (Topological Decomposition).
The 10-TET Tonnetz for the solution $(t=4,s=3)$ is isomorphic to the Levi graph of a $(10_{2})$ configuration, which decomposes into two disjoint Levi graphs of $(5_{2})$ configurations (cycles of length 10).
This means the harmonic Universe of $(4,3)$ solution 10-TET consists of two parallel "harmonic realities" that never intersect. Geometrically, the second graph is a perfect copy of the first, rotated by an angle of $\pi/10$ (one step in the circular arrangement). We display the separated components below, preserving their original geometric positions from Figure 6 to highlight the rotation.
Component A (The Blue Cycle)
$D_{0}$$M_{4}$$D_{4}$$M_{8}$$D_{8}$$M_{2}$$D_{2}$$M_{6}$$D_{6}$$M_{0}$
Component B (The Red Cycle)
$D_{1}$$M_{5}$$D_{5}$$M_{9}$$D_{9}$$M_{3}$$D_{3}$$M_{7}$$D_{7}$$M_{1}$
Figure 7. The decomposition of the 10-TET Tonnetz into two independent $(5_{2})$ Levi graphs. Note that the geometric shape of the red cycle is identical to the blue cycle but rotated by $18^{\circ}$, occupying the gaps left by the blue component.
7. The Functional Perspective: Restoring Connectivity in 3D
Our analysis of “modal degeneracy” in Theorem 1 relied on a strict set-theoretical perspective: since the pitch-class sets of $D_{r}$ and $M_{r+7}$ are identical, we treated them as the same vertex in the Tonnetz. This reduced the graph valence to 2 and caused the topological splitting described in Section 7.
However, from a functional (Riemannian) perspective, a chord is defined not just by its pitch content, but by its root. In this view, the major triad $D_{0}=(0,4,7)$ rooted at 0 is distinct from the minor triad $M_{7}=(7,0,4)$ rooted at 7, despite comprising the same notes. The “Relative” transformation ($R$) is no longer an identity map, but a bijection – a wormhole – mapping a chord from the “even universe” to the “odd universe” (since $r$ and $r+7$ have opposite parity).
If we adopt this functional differentiation, the vertex set doubles from 10 unique sets to 20 functional chords. The topology undergoes a dimensional lift:
  1. (1)
    The “Parallel” ($P$) and “Leading-Tone” ($L$) transformations continue to link chords within the same parity cycle (forming the Blue and Red rings derived in Section 7).
  2. (2)
    The “Relative” ($R$) transformation now acts as a transversal edge, connecting every node $D_{r}$ in the Blue cycle to a node $M_{r+7}$ in the Red cycle (and vice versa).
Consequently, the graph valence is restored to 3 (as in 12-TET), and the disconnected components merge into a single, connected 3-dimensional structure. The two disjoint cycles of the 2D projection become the bases of a **prism graph** (or cylindrical lattice) in 3D space.
We visualize this “Functional 10-TET Cylinder” below. The Blue and Red cycles form the top and bottom bases. To enhance readability and avoid visual overlapping, we apply a rotational offset to the Upper (Red) ring, creating a “twisted” cylinder where the functional $R$-connections form inclined bridges.
$D_{0}$$M_{4}$$D_{4}$$M_{8}$$D_{8}$$M_{2}$$D_{2}$$M_{6}$$D_{6}$$M_{0}$$M_{7}$$D_{7}$$M_{1}$$D_{1}$$M_{5}$$D_{5}$$M_{9}$$D_{9}$$M_{3}$$D_{3}$Even Universe (Sets)Odd Universe (Sets)Functional ConnectionThe R-Transformation
Figure 8. The “Functional Cylinder” of 10-TET. The Blue and Red cycles form the geometric bases, visualized here with alternating edge lengths to reflect the distinction between Parallel ($P$) and Leading-Tone ($L$) transformations. The gray dashed lines represent the Relative ($R$) transformation, acting as a wormhole between the disjoint pitch-class set universes.
7.1. Musical Implications: The Phantom Modulation
This functional restoration of connectivity in 3D has profound musical implications, extending the concept of “Janus-faced” triads introduced in Section 6.3.
In classical 12-TET harmony, modulation typically involves changing pitch classes (e.g., introducing a sharp or flat). However, in the $(4,3)$ solution 10-TET, the $R$-transformation represented by the vertical edges of the cylinder corresponds to a phantom modulation. The physical sound remains identical (the pitch-class set is invariant), but the listener is instantaneously transported from the “Even Universe” to the “Odd Universe” purely by re-interpreting the root.
This suggests that the $(4,3)$ solution 10-TET separates physical motion (changing notes) from functional motion (changing keys) more radically than 12-TET. Traversing the blue or red rings corresponds to physical changes in pitch, while traversing the vertical “wormholes” corresponds to purely cognitive shifts in harmonic function. The cylinder acts as a topological map of this duality, unifying the disjoint set-theoretical cycles into a single functional object.
7.2. The Combinatorial Structure: The $(10_{3})$ Configuration
By explicitly distinguishing between $D_{r}$ and $M_{r+7}$ and treating the identity relation as a valid edge ($R$), the system reveals a highly regular internal geometry. The interactions of $P$, $L$, and $R$ generate a unified graph that can be analyzed using the language of combinatorial design.
7.2.1. The Geometric Structure
In this restored view, every Major chord $D_{r}$ connects to exactly three Minor chords:
  • $M_{r}$ via the Parallel operation $P$ (intra-universe),
  • $M_{r+4}$ via the Leading-tone operation $L$ (intra-universe),
  • $M_{r+7}$ via the Wormhole operation $R$ (inter-universe).
Symmetrically, every Minor chord connects to exactly three Major chords. This transforms our graph into a connected, cubic (3-valent) bipartite graph with $V=20$ vertices ($10$ Majors and $10$ Minors).
This structure is a realization of a combinatorial $(10_{3})$ configuration. We can interpret the 10 Minor chords as “Points” and the 10 Major chords as “Lines”. The incidence relation is defined by our harmonic connections: each Line (Major chord) passes through exactly 3 Points (the Minors it connects to), and each Point (Minor chord) lies on exactly 3 Lines.
7.2.2. Visualization: The Levi Graphs
The incidence graph of a configuration is formally known as a Levi graph. In Figure 9, we present this graph in a layout corresponding to the spectral analysis from the previous sections. The Blue and Red universes form two concentric rings, with the $P$ and $L$ operations creating the perimeter of each ring. The $R$ operations (gold dashed lines) act as radial wormholes fusing the two worlds together.
$D_{9}$$M_{3}$$D_{3}$$M_{7}$$D_{7}$$M_{1}$$D_{1}$$M_{5}$$D_{5}$$M_{9}$$D_{6}$$M_{0}$$D_{0}$$M_{4}$$D_{4}$$M_{8}$$D_{8}$$M_{2}$$D_{2}$$M_{6}$Blue UniverseRed Universe
Figure 9. The Levi graph representing the $(10_{3})$ configuration. The Blue and Red universes are visualized as concentric rings, connected by radial $R$-wormholes (gold dashed lines), uniting the system.
In Figure 10, we present the same graph in the standard sequential notation, where chords are arranged $D_{0},M_{0},D_{1},M_{1}\dots$ along a single circle. This layout highlights the symmetry of the configuration and clearly shows the alternating nature of the Blue and Red subnetworks.
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$
Figure 10. The Connected Levi Graph in standard circular notation. Blue and Red nodes/edges denote the two universes; dashed gold lines denote the Identity Wormholes.
7.2.3. Minimal Chiral Cycles
The introduction of wormholes significantly reduces the girth (shortest cycle length) of the graph. While the intra-universe graphs have a girth of 10, the restored $(10_{3})$ graph allows for tight loops of length 4.
Theorem 3.
In the restored $(4,3)$ system, there are exactly two isomorphism classes of minimal cycles (of length 4) under the rotational symmetry of $\mathbb{Z}_{10}$. These classes are distinguished by their chirality (direction of winding through the wormhole).
Proof.
A cycle must alternate between Major and Minor chords. A length-2 cycle implies a double edge, which is impossible as $P,L,R$ are distinct operations. Thus, the minimal length is 4. A 4-cycle must involve at least two wormhole crossings to leave and return to the starting universe. Algebraically, we look for a sequence of operations that results in the identity element. The two minimal solutions correspond to the two directions of traversing the wormhole:
  1. (1)
    The Violet Cycle (Right-handed):
    $S=P\cdot R_{+7}\cdot L_{-4}\cdot R_{+7}$
    Tracing the index change: $P(0)\to R(+7)\to L(-4)\to R(+7)$. The total displacement is $7-4+7=+10\equiv 0\pmod{10}$.
  2. (2)
    The Orange Cycle (Left-handed):
    $S^{\prime}=P\cdot R_{+3}\cdot L_{+4}\cdot R_{+3}$
    Note that the inverse of the wormhole $D\to M_{+7}$ is $M\to D_{-7}=D_{+3}$. Tracing the index change: $P(0)\to R(+3)\to L(+4)\to R(+3)$. The total displacement is $3+4+3=+10\equiv 0\pmod{10}$.
These two sequences are non-isomorphic because they traverse the graph with opposite chirality, effectively winding around the toroidal structure of the configuration in opposite directions. ∎
Figure 11 illustrates these two fundamental cycles embedded within the Levi graph, depicted with straight edges and directional arrows to emphasize the geometric path.
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$$M_{0}$$D_{0}$$M_{7}$$D_{3}$$D_{5}$$M_{5}$$D_{8}$$M_{2}$
Violet Cycle ($S$): $M_{0}\to D_{0}\to M_{7}\to D_{3}$
Orange Cycle ($S^{\prime}$): $D_{5}\to M_{5}\to D_{8}\to M_{2}$
Figure 11. Visualisation of the two chiral minimal cycles. Straight edges connect the nodes, with arrows placed at the midpoint of each segment to indicate direction. These cycles represent the fundamental loops of the $(10_{3})$ configuration.
7.3. Algebraic Characterization and Graph Invariants
To properly classify the functional 10-TET graph within the family of cubic graphs, we must first establish its fundamental algebraic invariants. By inspection of the graph constructed in Section 7.2, we identify the following properties:
  1. (1)
    Order and Valence: The graph has $V=20$ vertices and is $3$-regular (cubic).
  2. (2)
    Bipartiteness: The graph is bipartite (Levi graph), with the two partitions corresponding to Major chords ($D$) and Minor chords ($M$).
  3. (3)
    Girth: As proven in Theorem 2, the graph has a girth of $g=4$ (containing 4-cycles such as $M_{0}\to D_{0}\to M_{7}\to D_{3}$).
  4. (4)
    Hamiltonicity: The graph is Hamiltonian. A complete tour of all 20 chords can be constructed by alternating the Wormhole ($R$) and Leading-Tone ($L$) operations.
7.3.1. The Grand Tour: A Hamiltonian Cycle
A specific Hamiltonian cycle can be generated by the operation sequence $(R\cdot L)^{10}$. Starting from $D_{0}$, this sequence visits every chord exactly once before returning to the start. The progression effectively transposes the chord by a minor third (interval class 3) at each $L$-step:
$D_{0}\xrightarrow{R}M_{7}\xrightarrow{L}D_{3}\xrightarrow{R}M_{0}\xrightarrow{L}D_{6}\xrightarrow{R}M_{3}\dots\xrightarrow{L}D_{0}$
Geometrically, this path zig-zags between the “Blue” and “Red” universes, weaving them into a single continuous thread.
$D_{9}$$M_{3}$$D_{3}$$M_{7}$$D_{7}$$M_{1}$$D_{1}$$M_{5}$$D_{5}$$M_{9}$$D_{6}$$M_{0}$$D_{0}$$M_{4}$$D_{4}$$M_{8}$$D_{8}$$M_{2}$$D_{2}$$M_{6}$$D_{0}$(a) Concentric Visualization
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$$D_{0}$(b) Circular Visualization
Figure 12. A Hamiltonian cycle ("The Grand Tour") visualized in two distinct layouts. The sequence $(RL)^{10}$ visits every chord exactly once. The path alternates between Identity Wormholes (dashed gold) and Leading-Tone steps (solid orange). (a) In the concentric layout, arrows indicate traversal order. (b) In the standard circular layout, the path forms a complex star-polygon.
7.4. The Classification Problem: Kantor, Martinetti, and the Missing Graph
Having established that our graph is a connected $(10_{3})$ configuration (in the combinatorial sense of being 3-regular with 10 “points” and 10 “lines”), we must check it against the historical classification of such structures.
In the late 19th century, Kantor (1881) [9] and Martinetti (1887) [10] rigorously classified the ten non-isomorphic connected $(10_{3})$ configurations. A comparison of invariants, however, yields a discrepancy: our functional 10-TET graph appears nowhere on this list.
This absence is not due to an error in the graph construction, nor an incompleteness in the classical list, but rather a divergence in the axiomatic definition of a “configuration.” The Kantor-Martinetti classification is grounded in projective geometry, which imposes the Axiom of Incidence:
Two distinct points are incident with at most one common line.
In graph-theoretical terms, this axiom forbids the existence of “digons” or “squares” (cycles of length 4). If a graph contains a 4-cycle ($P_{1}-L_{1}-P_{2}-L_{2}-P_{1}$), it implies that lines $L_{1}$ and $L_{2}$ intersect twice (at $P_{1}$ and $P_{2}$), which is geometrically degenerate. Therefore, all ten classical configurations possess a girth $g\geq 6$.
Since our graph possesses a girth $g=4$ (as shown in Section 7.3), it violates the geometric axiom of incidence. It is, therefore, a “degenerate” configuration in the eyes of classical geometry, but a perfectly valid (and physically meaningful) object in topological graph theory.
7.5. Identification: The Generalized Petersen Graph
To identify the precise isomorphism class of our graph, we turn to the generalized classification of cubic graphs provided by the Petersen $GP(10,k)$ family [2, 5].
Theorem 4.
The functional 10-TET graph is isomorphic to the Decagonal Prism Graph, also classified as the Generalized Petersen Graph $GP(10,1)$.
Proof.
The Generalized Petersen Graph $GP(n,k)$ is defined by vertices $\{u_{0},\dots,u_{n-1}\}\cup\{v_{0},\dots,v_{n-1}\}$ and edges:
$(u_{i},u_{i+1}),\quad(u_{i},v_{i}),\quad(v_{i},v_{i+k})$
By inspection of the 10-TET graph properties:
  1. (1)
    We identify two disjoint 10-cycles formed by the $P$ and $L$ operations: the Blue Ring ($u_{i}$) and the Red Ring ($v_{i}$).
  2. (2)
    The Wormhole operation $R$ connects each node in the Blue Ring to a unique node in the Red Ring, establishing the “spokes” $(u_{i},v_{i})$.
  3. (3)
    Unlike the Desargues graph (which is $GP(10,3)$), our rings preserve local connectivity without skipping steps in the permutation cycle.
Thus, setting $n=10$ and $k=1$, we observe that the structure is a prism with bases connected by identity wormholes. This confirms the isomorphism to $GP(10,1)$. ∎
8. The “Wide” Choice of Thirds: A Geometric Revelation
We now return to the singular case identified in Section 5.4: the “Wide” 10-TET system defined by the parameters $(t,s)=(6,1)$. While the earlier sections focused on the topological continuity of the “narrow” systems (where $\Delta=1$), the $\Delta=5$ case offers a different kind of beauty. It is not merely another valid tuning; as we shall see, it turns out to be the only configuration in the 10-TET universe that realizes a fully vertex-transitive Levi graph structure.
In the following analysis, we demonstrate that this musical system is not just a graph of chords, but an exact musical isomorphism of the famous Desargues configuration $(10_{3})_{1}$. This result bridges the gap between harmonic theory and projective geometry, revealing that the most symmetric musical choice in 10-TET corresponds to one of the most fundamental theorems in mathematics.
8.1. Derivation
We solve the system of congruences:
(10) $\begin{cases}t+s\equiv 7\\ t-s\equiv 5\end{cases}\pmod{10}$
Summing the equations yields $2t\equiv 12\equiv 2\pmod{10}$, which implies $t=1$ or $t=6$. Since a major third of 1 step is counter-intuitive, we select:
$\mathbf{t=6},\quad\mathbf{s=1}.$
8.2. Definition of Chords
Using the solution $(6,1)$, we define the chords for every root $r\in\{0,\dots,9\}$.
  • Major Chords ($D_{r}$): Defined as $(r,r+t,r+q)=\{r,r+6,r+7\}$.
  • Minor Chords ($M_{r}$): Defined as $(r,r+s,r+q)=\{r,r+1,r+7\}$.
Table 4. Table of Major Chords ($D_{r}$) for $t=6$
Name Root Set
$D_{0}$ 0 {0, 6, 7}
$D_{1}$ 1 {1, 7, 8}
$D_{2}$ 2 {2, 8, 9}
$D_{3}$ 3 {3, 9, 0}
$D_{4}$ 4 {4, 0, 1}
$D_{5}$ 5 {5, 1, 2}
$D_{6}$ 6 {6, 2, 3}
$D_{7}$ 7 {7, 3, 4}
$D_{8}$ 8 {8, 4, 5}
$D_{9}$ 9 {9, 5, 6}
Table 5. Table of Minor Chords ($M_{r}$) for $s=1$
Name Root Set
$M_{0}$ 0 {0, 1, 7}
$M_{1}$ 1 {1, 2, 8}
$M_{2}$ 2 {2, 3, 9}
$M_{3}$ 3 {3, 4, 0}
$M_{4}$ 4 {4, 5, 1}
$M_{5}$ 5 {5, 6, 2}
$M_{6}$ 6 {6, 7, 3}
$M_{7}$ 7 {7, 8, 4}
$M_{8}$ 8 {8, 9, 5}
$M_{9}$ 9 {9, 0, 6}
The interval vectors for Major $[6,1,3]$ and Minor $[1,6,3]$ are distinct. Thus, **no modal degeneracy exists**.
8.3. Connectivity and the Tonnetz
We construct the Tonnetz by connecting chords that share exactly two common tones. The system is perfectly 3-valent.
  • P
    (Parallel): $D_{r}\leftrightarrow M_{r}$ (Share root and fifth).
  • L
    (Leading): $D_{r}\leftrightarrow M_{r+6}$ (Share third and fifth).
  • R
    (Relative): $D_{r}\leftrightarrow M_{r+9}$ (Share root and third).
$D_{r}$ Connections Chord Neighbors $\{M_{i},M_{j},M_{k}\}$ $D_{0}$ $M_{0},M_{6},M_{9}$ $D_{1}$ $M_{1},M_{7},M_{0}$ $D_{2}$ $M_{2},M_{8},M_{1}$ $D_{3}$ $M_{3},M_{9},M_{2}$ $D_{4}$ $M_{4},M_{0},M_{3}$ $D_{5}$ $M_{5},M_{1},M_{4}$ $D_{6}$ $M_{6},M_{2},M_{5}$ $D_{7}$ $M_{7},M_{3},M_{6}$ $D_{8}$ $M_{8},M_{4},M_{7}$ $D_{9}$ $M_{9},M_{5},M_{8}$
$M_{r}$ Connections Chord Neighbors $\{D_{i},D_{j},D_{k}\}$ $M_{0}$ $D_{0},D_{4},D_{1}$ $M_{1}$ $D_{1},D_{5},D_{2}$ $M_{2}$ $D_{2},D_{6},D_{3}$ $M_{3}$ $D_{3},D_{7},D_{4}$ $M_{4}$ $D_{4},D_{8},D_{5}$ $M_{5}$ $D_{5},D_{9},D_{6}$ $M_{6}$ $D_{6},D_{0},D_{7}$ $M_{7}$ $D_{7},D_{1},D_{8}$ $M_{8}$ $D_{8},D_{2},D_{9}$ $M_{9}$ $D_{9},D_{3},D_{0}$
Figure 13. Adjacency tables for the Wide System ($\Delta=5$). Note that the graph mixes indices differently than the compact system, ensuring global connectivity.
8.4. The Levi Graph
The resulting graph connects all 20 chords into a single structure. We visualize it by placing chords on a circle: $D_{0},M_{0},D_{1},M_{1}\dots$. The $P$ and $R$ transformations form the perimeter, while $L$ forms the internal chords.
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$Levi Graph ($\Delta=5$)
Figure 14. The Levi Graph of the Wide System ($\Delta=5$). All 20 chords form a single connected component.
8.5. Topology of Cycles: Rigorous Proof of Girth 6
A fundamental question in the topology of this harmonic network is the length of the shortest path to return to the starting chord (the **girth**). Here we present a rigorous proof using the "Shift Method".
8.6. Parity Constraint
Our harmonic graph is **bipartite**: edges only connect Major ($D$) chords to Minor ($M$) chords. Consequently, any walk starting at a Major chord will be at a Minor chord after an odd number of steps (1, 3, 5…). To return to the starting type (Major), the path length must necessarily be **even**. Therefore, cycles of length 3 or 5 cannot exist. We need only examine lengths $2k$ (i.e., 2, 4, 6…).
8.7. The "Shift Method" for Even Cycles
We analyze the path as a sequence of "double-steps": $D_{start}\to M\to D_{next}$. Each double-step shifts the root index by a specific amount modulo 10. To close a cycle, the sum of these shifts must be $0\equiv 10\pmod{10}$.
There are exactly 6 possible distinct double-steps starting from $D_{r}$:
  1. (1)
    PR (via $M_{r}$): $D_{r}\to M_{r}\to D_{r+1}$. Shift: $\mathbf{+1}$.
  2. (2)
    PL (via $M_{r}$): $D_{r}\to M_{r}\to D_{r-6}$. Shift: $\mathbf{+4}$ (since $6\equiv 4$).
  3. (3)
    LP (via $M_{r+6}$): $D_{r}\to M_{r+6}\to D_{r+6}$. Shift: $\mathbf{+6}$.
  4. (4)
    LR (via $M_{r+6}$): $D_{r}\to M_{r+6}\to D_{r+7}$. Shift: $\mathbf{+7}$.
  5. (5)
    RP (via $M_{r-1}$): $D_{r}\to M_{r-1}\to D_{r-1}$. Shift: $\mathbf{+9}$ (inverse of PR).
  6. (6)
    RL (via $M_{r-1}$): $D_{r}\to M_{r-1}\to D_{r-7}$. Shift: $\mathbf{+3}$ (inverse of LR).
8.8. Proof that Length 4 is Impossible
A cycle of length 4 consists of exactly **two** double-steps. For a non-trivial cycle (one that does not simply backtrack like $PR$ then $RP$), we must select two different non-inverse shifts that sum to $10$. Let us examine the possible sums of distinct shifts from our set $\{1,3,4,6,7,9\}$:
  • $1+3=4$, $1+4=5$, $1+6=7$, $1+7=8$.
  • $3+4=7$, $3+6=9$.
  • $4+6=10$. This looks promising! However, shift +4+4 is $PL$ and shift +6+6 is $LP$. The sequence would be $PL$ followed by $LP$, i.e., $P\to L\to L\to P$. Since $L^{2}=Identity$, this collapses to $P\to P=Identity$. This is a backtracking path, not a cycle.
  • The same logic applies to $1+9=10$ (PR and RP) and $3+7=10$ (RL and LR).
Since no two distinct, non-inverse shifts sum to 10, **cycles of length 4 do not exist.**
8.9. Proof that Length 6 Exists
A cycle of length 6 requires **three** double-steps. We look for three shifts that sum to 10. We find the solution:
$1+6+3=10$
Mapping these numbers back to operations:
  • Shift $\mathbf{+1}$: Operation $PR$.
  • Shift $\mathbf{+6}$: Operation $LP$.
  • Shift $\mathbf{+3}$: Operation $RL$.
Linking these sequentially creates the path: $(PR)\to(LP)\to(RL)$. This corresponds to the sequence $P,R,L,P,R,L$. This is exactly the $(PRL)^{2}$ cycle.
8.10. Chirality
Because the operations do not commute, the order matters.
  • **Right-Handed Cycle $(PRL)^{2}$:** Uses the sum $1+6+3$.
  • **Left-Handed Cycle $(PLR)^{2}$:** Corresponds to the shifts $4+9+7=20\equiv 0$.
These are the two shortest paths through the harmonic space, visualized below.
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$Cycle $(PLR)^{2}$Cycle $(PRL)^{2}$
Figure 15. Visualization of the two chiral shortest cycles (girth 6). To avoid visual overlap, we present the Left-Handed cycle (Violet) starting at $D_{0}$, and the Right-Handed cycle (Orange) starting at $D_{8}$. Arrows indicate the direction of the flow, revealing the distinct chiral nature of each path.
8.11. Geometric Identification: The Desargues Configuration
The structural properties of the derived Tonnetz lead to a significant geometric identification. We have established that the graph is:
  1. (1)
    **Cubic (3-valent):** Every chord has exactly 3 neighbors.
  2. (2)
    **Bipartite:** Connections only exist between Major and Minor chords.
  3. (3)
    **Order 20:** There are 10 Major and 10 Minor chords.
  4. (4)
    **Girth 6:** The shortest cycle has length 6.
  5. (5)
    **Vertex-Transitive:** Due to the cyclic generation of chords via $\mathbb{Z}_{10}$ and the uniform definition of transformations $P,L,R$, the graph exhibits a high degree of symmetry. Mapping $D_{r}\to D_{r+k}$ implies a corresponding automorphism of the entire graph.
In the classification of combinatorial configurations, a system of $n$ points and $n$ lines where each point lies on 3 lines and each line contains 3 points is denoted as a $(n_{3})$ configuration. The Levi graph of such a configuration connects points to lines, resulting in a bipartite cubic graph of $2n$ vertices.
For $n=10$, there are exactly 10 distinct combinatorial types of $(10_{3})$ configurations (see [1] Table 1, p.333, or [14], Section 5.3.4, p. 176 - 179). Among these, **only one** possesses a vertex-transitive Levi graph: the **Desargues Configuration**, usually denoted as $(10_{3})_{1}$.
Theorem 5.
The Levi graph of the Wide 10-TET system ($\Delta=5$) is isomorphic to the **Desargues Graph**.
Proof.
The Desargues graph is the unique symmetric cubic graph on 20 vertices. Since our harmonic graph is generated by a vertex-transitive group action on $\mathbb{Z}_{10}$ and matches the order and valency requirements, it must be isomorphic to the Desargues graph. ∎
This identification is profound. The Desargues configuration is the geometric embodiment of Desargues’ Theorem, a fundamental result in projective geometry. Unlike other configurations that may require specific fields, the Desargues configuration can be realized in any Desarguesian projective plane $PG(2,\mathbb{K})$, including the real projective plane and finite fields $\mathbb{K}$ such as $\mathbb{F}_{5}$.
Thus, the "Wide" 10-TET harmonic system is not merely a graph of chords; it is a musical realization of one of the most classic structures in projective geometry. The Major chords can be viewed as "Points" and the Minor chords as "Lines" (or vice versa) in a projective configuration where harmonic relations define incidence.
9. The Tritone Solution: System (5,2)
In Section 6.3, we briefly encountered the “neutral” solution to the Pythagorean constraints, characterized by the difference $\Delta=3$.
(11) $\begin{cases}t+s\equiv 7\\ t-s\equiv 3\end{cases}\implies 2t\equiv 10\pmod{10}$
This yields the solution $t=5$ and $s=2$. Although we initially set this aside due to the acoustic strangeness of treating the tritone ($t=5$) as a consonance, the topological robustness of this system merits a full analysis. As we shall see, this system serves as a fascinating “middle ground” between the disconnected Acoustic system ($\Delta=1$) and the highly symmetric Wide system ($\Delta=5$).
9.1. Chord Definitions and Chirality
We define the chords with root $r$ using the intervals $t=5$ (Major Third / Tritone) and $s=2$ (Minor Third / Whole Tone):
(12) $D_{r}$ $=\{r,r+5,r+7\}\pmod{10}$
(13) $M_{r}$ $=\{r,r+2,r+7\}\pmod{10}$
A unique feature of this system is the intervallic symmetry of the chords. For a Major chord $D_{0}=\{0,5,7\}$, the interval vector is $(5,2,3)$. For a Minor chord $M_{0}=\{0,2,7\}$, the interval vector is $(2,5,3)$. Geometrically, these triads are mirror images (chiral reflections) of each other within the chromatic circle, sharing the perfect fifth $\{0,7\}$ as a rigid spine.
The complete list of chords generated by this system is provided in Tables 7 and 7.
Table 6. Major Chords $D_{r}=\{r,r+5,r+7\}$
Chord Pitch Class Set
$D_{0}$ $\{0,5,7\}$
$D_{1}$ $\{1,6,8\}$
$D_{2}$ $\{2,7,9\}$
$D_{3}$ $\{3,8,0\}$
$D_{4}$ $\{4,9,1\}$
$D_{5}$ $\{5,0,2\}$
$D_{6}$ $\{6,1,3\}$
$D_{7}$ $\{7,2,4\}$
$D_{8}$ $\{8,3,5\}$
$D_{9}$ $\{9,4,6\}$
Table 7. Minor Chords $M_{r}=\{r,r+2,r+7\}$
Chord Pitch Class Set
$M_{0}$ $\{0,2,7\}$
$M_{1}$ $\{1,3,8\}$
$M_{2}$ $\{2,4,9\}$
$M_{3}$ $\{3,5,0\}$
$M_{4}$ $\{4,6,1\}$
$M_{5}$ $\{5,7,2\}$
$M_{6}$ $\{6,8,3\}$
$M_{7}$ $\{7,9,4\}$
$M_{8}$ $\{8,0,5\}$
$M_{9}$ $\{9,1,6\}$
9.2. Adjacency and Graph Connectivity
We determine the Tonnetz connectivity by identifying neighbors that share exactly two common tones.
9.2.1. The P-Transformation (Parallel)
$D_{r}$ shares $\{r,r+7\}$ (Root, Fifth) with a minor chord. Looking at $M_{r}=\{r,r+2,r+7\}$, we see an immediate match.
$D_{r}\leftrightarrow M_{r}$
This operation preserves the root and fifth, moving the third by $\Delta=3$ steps.
9.2.2. The R-Transformation (Relative)
$D_{r}$ shares $\{r,r+5\}$ (Root, Third) with a minor chord. We seek $M_{k}=\{k,k+2,k+7\}$ containing $\{r,r+5\}$. Setting $k=r+8$ (which is r2r-2), we get $M_{r+8}=\{r+8,r,r+5\}$.
$D_{r}\leftrightarrow M_{r+8}$
This operation connects the Major chord at index $r$ to the Minor chord shifted by 2-2 steps.
9.2.3. The L-Transformation (Leading-Tone)
$D_{r}$ shares $\{r+5,r+7\}$ (Third, Fifth) with a minor chord. We seek $M_{k}$ containing these tones. Setting $k=r+5$, we get $M_{r+5}=\{r+5,r+7,r+2\}$.
$D_{r}\leftrightarrow M_{r+5}$
This is a tritone jump. The $L$ transformation connects a Major chord to the Minor chord rooted at its tritone antipode.
The full set of connections for the system is given in Tables 9 and 9.
Table 8. Major Chord Connections ($D_{r}\to M_{k}$)
Chord P ($M_{r}$) L ($M_{r+5}$) R ($M_{r-2}$)
$D_{0}$ $M_{0}$ $M_{5}$ $M_{8}$
$D_{1}$ $M_{1}$ $M_{6}$ $M_{9}$
$D_{2}$ $M_{2}$ $M_{7}$ $M_{0}$
$D_{3}$ $M_{3}$ $M_{8}$ $M_{1}$
$D_{4}$ $M_{4}$ $M_{9}$ $M_{2}$
$D_{5}$ $M_{5}$ $M_{0}$ $M_{3}$
$D_{6}$ $M_{6}$ $M_{1}$ $M_{4}$
$D_{7}$ $M_{7}$ $M_{2}$ $M_{5}$
$D_{8}$ $M_{8}$ $M_{3}$ $M_{6}$
$D_{9}$ $M_{9}$ $M_{4}$ $M_{7}$
Table 9. Minor Chord Connections ($M_{r}\to D_{k}$)
Chord P ($D_{r}$) L ($D_{r+5}$) R ($D_{r+2}$)
$M_{0}$ $D_{0}$ $D_{5}$ $D_{2}$
$M_{1}$ $D_{1}$ $D_{6}$ $D_{3}$
$M_{2}$ $D_{2}$ $D_{7}$ $D_{4}$
$M_{3}$ $D_{3}$ $D_{8}$ $D_{5}$
$M_{4}$ $D_{4}$ $D_{9}$ $D_{6}$
$M_{5}$ $D_{5}$ $D_{0}$ $D_{7}$
$M_{6}$ $D_{6}$ $D_{1}$ $D_{8}$
$M_{7}$ $D_{7}$ $D_{2}$ $D_{9}$
$M_{8}$ $D_{8}$ $D_{3}$ $D_{0}$
$M_{9}$ $D_{9}$ $D_{4}$ $D_{1}$
9.3. The Levi Graph
Since the chord sets $D$ and $M$ are disjoint, the resulting structure is a cubic bipartite graph of order 20. The $P$ and $R$ transformations preserve the parity of the root index ($D_{2k}\leftrightarrow M_{2k}$ and $D_{2k}\leftrightarrow M_{2k-2}$), creating two disjoint cycles of 10 chords each. These are the “Even Universe” (Blue) and the “Odd Universe” (Red). The system becomes fully connected via the $L$ transformation, which maps $D_{r}\leftrightarrow M_{r+5}$, effectively linking an even root to an odd root.
The visualization of this graph is presented in Figure 16.
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$Levi Graph of (5,2)
Figure 16. The Levi graph of the (5,2) system arranged on a single circle. Nodes are colored by parity (Blue=Even, Red=Odd). The connections show the P (adjacent), R (short chord), and L (long dashed chord) transformations. Note that L connects the blue and red sub-universes.
9.4. Minimal Cycles
To characterize the local topology of the graph, we analyze the minimal cycles (girth). Unlike the $(4,3)$ system which possessed a hexagonal structure ($g=6$), the $(5,2)$ system exhibits a tighter connectivity.
Theorem 6.
The Tritone graph has a girth of $g=4$. The minimal cycles are of the form $(PL)^{2}$, forming squares that connect antipodal chords.
Proof.
Consider the sequence of operations starting from $D_{r}$:
  1. (1)
    $D_{r}\xrightarrow{P}M_{r}$.
  2. (2)
    $M_{r}\xrightarrow{L}D_{r+5}$.
  3. (3)
    $D_{r+5}\xrightarrow{P}M_{r+5}$.
  4. (4)
    $M_{r+5}\xrightarrow{L}D_{r+5+5}=D_{r+10}=D_{r}$.
This sequence $D_{r}\to M_{r}\to D_{r+5}\to M_{r+5}\to D_{r}$ forms a closed cycle of length 4. Since the graph is bipartite, it contains no odd cycles (length 3). Therefore, $g=4$. ∎
Geometrically, these cycles appear as “bow-tie” structures crossing the center of the circular Levi graph, linking an Even pair ($D_{r},M_{r}$) with its antipodal Odd pair ($D_{r+5},M_{r+5}$).
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$$D_{0}$$M_{0}$$D_{5}$$M_{5}$$D_{7}$$M_{7}$$D_{2}$$M_{2}$
Figure 17. Two minimal cycles (girth $g=4$) of opposite chirality. The Violet path ($D_{0}\to M_{0}\to D_{5}\to M_{5}$) traverses locally clockwise, while the Orange path ($M_{2}\to D_{2}\to M_{7}\to D_{7}$) traverses locally counter-clockwise.
9.4.1. Algebraic Characterization
We summarize the graph invariants for the (5,2) solution:
  1. (1)
    Order and Valence: The graph has $V=20$ vertices and is $3$-regular (cubic).
  2. (2)
    Bipartiteness: The graph is bipartite (Levi graph).
  3. (3)
    Girth: The girth is $g=4$, characterized by the $(PL)^{2}$ squares.
  4. (4)
    Hamiltonicity: The graph is Hamiltonian, as proved by the construction of the “Tritone Tour” in the following section.
9.5. The Tritone Tour
Because the system is connected, finding a Hamiltonian cycle is straightforward. We can traverse the “Even Universe” almost entirely, cross via a tritone bridge ($L$) to the “Odd Universe”, traverse it, and bridge back.
A valid Hamiltonian sequence is:
$(P\cdot R)^{4}\cdot L\cdot(R\cdot P)^{4}\cdot R\cdot L\cdot R$
The Path Trace:
  1. (1)
    Even Sector: $D_{0}\to M_{0}\to D_{2}\to M_{2}\to D_{4}\to M_{4}\to D_{6}\to M_{6}\to D_{8}$.
  2. (2)
    Bridge: $D_{8}\xrightarrow{L}M_{3}$ (Switch to Odd).
  3. (3)
    Odd Sector: $M_{3}\to D_{5}\to M_{5}\to D_{7}\to M_{7}\to D_{9}\to M_{9}\to D_{1}\to M_{1}$.
  4. (4)
    Return: $M_{1}\xrightarrow{R}D_{3}\xrightarrow{L}M_{8}\xrightarrow{R}D_{0}$.
$D_{0}$$M_{0}$$D_{2}$$M_{2}$$D_{4}$$M_{4}$$D_{6}$$M_{6}$$D_{8}$$M_{8}$$D_{1}$$M_{1}$$D_{3}$$M_{3}$$D_{5}$$M_{5}$$D_{7}$$M_{7}$$D_{9}$$M_{9}$
Figure 18. The Hamiltonian "Tritone Tour" in the (5,2) system. The path traverses the blue even ring, jumps via an orange tritone bridge ($L$) to the red odd ring, and returns to close the cycle.
9.6. Identification with the Generalized Petersen Graph
Having analyzed the local and global properties of the $(5,2)$ system, we can now formally identify its underlying graph structure.
Theorem 7.
The Levi graph of the (5,2) system is isomorphic to the Decagonal Prism Graph, which corresponds to the Generalized Petersen Graph $GP(10,1)$.
Proof.
The Generalized Petersen Graph $GP(n,k)$ consists of vertices $\{u_{0},\dots,u_{n-1}\}\cup\{v_{0},\dots,v_{n-1}\}$ with edges:
  1. (1)
    $u_{i}\sim u_{i+1}$ (Outer rim),
  2. (2)
    $u_{i}\sim v_{i}$ (Spokes),
  3. (3)
    $v_{i}\sim v_{i+k}$ (Inner rim).
In our system, we identified two disjoint 10-cycles generated by the $P$ and $R$ transformations: the Even Universe and the Odd Universe. Let the vertices of the Even Universe be mapped to the outer rim $u_{i}$, and the vertices of the Odd Universe be mapped to the inner rim $v_{i}$, such that the $L$-transformation corresponds to the spokes connecting them.
Specifically, the $L$-transformation connects $D_{r}$ (Even) to $M_{r+5}$ (Odd). If we align the indices such that the $L$-connection is vertical ($u_{i}\sim v_{i}$), the connectivity within the Odd universe (inner rim) mirrors the connectivity of the Even universe (outer rim) with a step size of $k=1$. Thus, the graph is a prism graph $Y_{10}$, which is denoted as $GP(10,1)$. ∎
10. Isomorphism of the Acoustic and Tritone Systems
We have arrived at a remarkable theoretical unification. In Section 7, we established that the “Restored” Acoustic system (4,3)—formed by connecting two disjoint chromatic universes via “wormhole” $R$-transformations—possesses the topology of the Decagonal Prism $GP(10,1)$. Similarly, in Section 9, we proved that the Tritone system (5,2)—naturally connected by $L$-transformations—also forms the Decagonal Prism $GP(10,1)$.
This implies that despite their radically different intervallic generators ($t=4,s=3$ versus $t=5,s=2$) and different consonance criteria, these two harmonic systems are isomorphic. They describe the exact same pathway through the space of triads, merely labelled with different pitch-classes.
10.1. The Isomorphism Theorem
To formalize this equivalence, we construct a mapping $\Phi$ that transforms the chord space of the Tritone system into the chord space of the Acoustic system. The key insight is that the topological role of the transformations is swapped:
  • The Rim of the prism in (5,2) is generated by $P$ and $R$. In (4,3), the Rim is generated by $P$ and $L$.
  • The Spokes of the prism in (5,2) are the $L$ (tritone) steps. In (4,3), the Spokes are the $R$ (wormhole) steps.
Theorem 8 (Isomorphism of 10-TET Systems).
Let $\mathcal{G}_{(5,2)}$ be the Levi graph of the Tritone system and $\mathcal{G}_{(4,3)}$ be the Levi graph of the Restored Acoustic system. There exists a graph isomorphism $\Phi:\mathcal{G}_{(5,2)}\to\mathcal{G}_{(4,3)}$ defined by the index mapping $\sigma(k)=7k\pmod{10}$ acting on the chord roots, such that:
(14) $\Phi(D_{k}^{(5,2)})$ $=D_{\sigma(k)}^{(4,3)}$
(15) $\Phi(M_{k}^{(5,2)})$ $=M_{\sigma(k)}^{(4,3)}$
This mapping induces a transformation of the harmonic operations:
(16) $\Phi(P_{(5,2)})=P_{(4,3)},\quad\Phi(R_{(5,2)})=L_{(4,3)},\quad\Phi(L_{(5,2)})=R_{(4,3)}.$
Proof.
The cycle of roots in the Even Universe of the (5,2) system proceeds by steps of +2+2 (via the $R$ operation). The cycle of roots in the First Universe of the (4,3) system proceeds by steps of +4+4 (via the $L$ operation). The mapping $\sigma(k)=7k$ maps the step generator $2$ to $4$, since:
$\sigma(k+2)=7(k+2)=7k+14\equiv 7k+4\pmod{10}$
Thus, adjacency along the rims is preserved. Similarly, the “Spoke” operation $L$ in (5,2) connects $D_{k}$ to $M_{k+5}$. The mapping $\sigma$ transforms this to a connection between $D_{7k}$ and $M_{7(k+5)}=M_{7k+35}\equiv M_{7k+5}$. In the (4,3) system, the $R$ operation connects chords with interval difference 5 (antipodal). Thus, the spokes are also mapped to spokes. Since $\Phi$ is a bijection preserving all edges, the graphs are isomorphic. ∎
10.2. The Double Grand Tour
The existence of this isomorphism allows us to translate any musical progression from one system to the other. In Figure 19, we illustrate the Hamiltonian Cycle found in Section 9 (The Tritone Tour) and its direct image under $\Phi$ in the Acoustic System.
For visual clarity, both graphs are presented on the standard circular Levi arrangement ($D_{0},M_{0},D_{1},M_{1}\dots$). The path on the left is the “Tritone Tour”. The path on the right is the isomorphic “Acoustic Tour”. The correspondence is indicated by the step numbers: the transition labeled $k$ on the left maps to the transition $\Phi(k)$ on the right.
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$1234567891011121314151617181920(a) System (5,2)
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$$\Phi_{1}$$\Phi_{2}$$\Phi_{3}$$\Phi_{4}$$\Phi_{5}$$\Phi_{6}$$\Phi_{7}$$\Phi_{8}$$\Phi_{9}$$\Phi_{10}$$\Phi_{11}$$\Phi_{12}$$\Phi_{13}$$\Phi_{14}$$\Phi_{15}$$\Phi_{16}$$\Phi_{17}$$\Phi_{18}$$\Phi_{19}$$\Phi_{20}$(b) System (4,3)
Figure 19. Visualizing the Parallel Grand Tours. Both systems are represented on the standard chromatic circle. Left: The Hamiltonian cycle in the (5,2) system. Right: The isomorphic cycle in the (4,3) system. The isomorphism $\Phi$ maps the Tritone Bridges (Orange, left) to the Wormhole Connections (Violet, right).
10.3. Musical Interpretation
The formal isomorphism reveals a profound duality between the two systems. In the Tritone System (5,2), the “difficult” interval is the Tritone ($t=5$). It is the interval that forces a jump between the Even and Odd parities. In the Acoustic System (4,3), the “difficult” interval is the Wormhole (antipodal connection). It is the interval that forces a jump between the two disjoint hexatonic-like universes.
Mathematically, Parity $\cong$ Universe. Navigating between Even and Odd chords using tritones in (5,2) is structurally indistinguishable from navigating between parallel harmonic universes using wormholes in (4,3). The “Grand Tour” of the 10-TET scale is thus a universal invariant: whether one views it through the lens of (5,2) or (4,3), the path requires exactly one bridge-crossing (and one return) to visit every harmony.
Acknowledgements
First and foremost, I dedicate this work to Leszek Możdżer, whose invitation to the decaphonic piano project sparked this entire mathematical journey. I am also deeply indebted to Lane P. Hughston, who made a special trip to Warsaw to introduce me to the theory of Tonnetze. I am profoundly grateful for his generosity in sharing his time to teach me the foundations of these networks and for illuminating their modern revival in the mathematics of music.
I am also deeply grateful to A. Bogucki, A. Włodarczyk, D. An, M. Demiański, C. D. Hill, M. Kuś, M. Mantegazza and R. Zduniak for their valuable discussions and support. Finally, special thanks are due to the Directors and Staff of the Center for Theoretical Physics of PAS for their patience and for tolerating such an unconventional musical project within the walls of a theoretical physics institute.
Appendix A Appendix A: Historical Motivation and the Acoustic Implementation
As written in the original paper on Pythagorean tunings for $n$-TETs [11], the author’s initial motivation for considering the 10-TET system comes from a live concert experience, summarized below.
In January 2023 I contacted the world renowned jazz pianist Leszek Możdżer, to ask if he would be willing to give a piano recital for the participants of a mathematics conference ‘GRIEG meets Chopin’ that I was helping to organize. To my surprise Leszek Możdżer’s answer to my request was positive, but with one caveat. Specifically, would I help him with the mathematics needed to redesign his Östlind and Almquist concert piano from the usual 12-step equally tempered (TET) scale to the 10-step TET scale? Możdżer further proposed that, at the concert during the mathematical conference, two pianos would be played: his redesigned 10-TET acoustic Östlind and Almquist piano and the usual 12-TET Steinway piano.
Without thinking much about ‘why the ten-step scale?’ I gladly accepted Możdżer’s proposal and in short order prepared a table with data needed to retune the piano from 12-TET to 10-TET. But while the mathematics I used was straightforward the actual retuning of the 12-TET piano to the 10-TET scale encountered many technical issues (the full technical procedure is described in [13]). These were eventually resolved by the combined efforts of two teams of Leszek Możdżer (with Roman Galiński, Jan Grzyśka, Ryszard Mariański, Mirosław Mastalerz, Sławomir Rosa) and mine (Aleksander Bogucki, Andrzej Włodarczyk). My team even filed a patent application with the major ideas of this retuning.
As a result the World Premiere of the Acoustic Decaphonic Piano by Leszek Możdżer took place on July 13, 2023 in the Nowa Miodowa Concert Hall in Warsaw, Poland. In a brilliant program he premiered a number of his compositions written for two pianos, both traditional 12-TET and the decaphonic 10-TET, some of which were played on both instruments simultaneously. He also performed a number of world’s piano masterpieces paraphrased for the 10-TET acoustic piano. In the opinion of many of those in attendance, he proved that with his virtuosity and for his musical compositions/paraphrases, the 10-TET piano is a wonderful instrument.
Leszek Możdżer’s answer to my question ‘why you want a 10-scale piano?’ is beyond the scope of this note; shortly: it was quite unsatisfactory for me. So, since January 2023 I have been looking for a mathematical argument that would characterize the 10-step musical scale among all other scales. The present paper is an attempt for such a characterization.